OpenStudy (anonymous):

W={(a1, a2) in R^2| a1+2a2=0} is W a subspace?

OpenStudy (anonymous):

It is the kernel of a linear map, hence it is a subspace.

OpenStudy (anonymous):

how do you prove that it is a subspace?

OpenStudy (anonymous):

\[ \mbox{Let}\;\;L:\mathbb{R}^2\rightarrow\mathbb{R}\;\mbox{with}\;L(x,y)=x+2y.\]

OpenStudy (anonymous):

Then L is a linear map of R-vectorspaces.

OpenStudy (anonymous):

Hence, \[ ker L:=\{ (x,y)\in\mathbb{R}^2\;|\;L(x,y)=0 \}=W \] is a subspace of |R^2.

OpenStudy (anonymous):

how do you prove that it is a subspace?

OpenStudy (anonymous):

I just proved it to you! The preimage of 0 under a linear map is a subspace. That is a theorem of linear algebra which you can use.

OpenStudy (anonymous):

In case you don't know that theorem, you may also prove the assertion directly.

OpenStudy (anonymous):

Would you like me to show you how to do it without use of linear maps?

OpenStudy (anonymous):

yes please, we haven't learned linear maps yet

OpenStudy (anonymous):

Ok, then take (x,y) in W and lambda in |R. Then (lambda*x, lambda*y) is also in W because if x+2y=0, then also 0=lambda*(x+2y)=lambda*x+2*(lambda*y). Now take (x,y) in W and (v,w) in W. Then (x+v, y+w) is also in W, because if x+2y=0 and v+2w=0, then also 0=(x+2y)+(v+2w)=(x+v)+2*(y+w). So, W is closed under scalar multiplication and vector addition and is thus a linear subspace of |R^2.